OpenStudy (anonymous):

Solve the inequality 1/(2-X) < 1

OpenStudy (anonymous):

1 < 2-X -1 < -X x < 1

OpenStudy (anonymous):

yea i shoulda wrote it out

OpenStudy (anonymous):

you are going to have to start with \[\frac{1}{2-x}-1<0\] and go from there

OpenStudy (anonymous):

@jimmyrep you cannot multiply by 2 - x because you don't know whether it is positive or negative

OpenStudy (anonymous):

Okay Thanks!

OpenStudy (anonymous):

it is true that \[x<1\] is a solution, but so is \[x>2\]

OpenStudy (anonymous):

you get \[\frac{x-1}{2-x}<0\] and check on the intervals \[(-\infty, 1), (1,2),(2, \infty)\] and see that first and third work

OpenStudy (anonymous):

i agree with what you said - you are right - its not valid to multiply by 2-x

OpenStudy (anonymous):

checking on the intervals is valid